Required information Camp Rainbow offers overnight summer camp programs for children ages 10–14 every summer during June andRequired 4 Required 5 Using the high-low method, calculate Camp Rainbows total fixed operating costs and variable operatingRequired 4 Required 5 Using the high-low method results, calculate the camps expected operating cost if 166 children attend
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity)/(Highest activity-Lowest activity)

= (14,000-8,060)/(194-106)

= 67.50

Fixed cost = Total cost - Total Variable cost

= 14,000 - (67.50*194)

= 905

Fixed cost 905
Variable cost per child 67.50

.

Expected cost

= (67.50*166) + 905

= 12,110